OpenStudy (anonymous):

-tan^2x+sec^2x= 1 ? match the right to the left. help ?

OpenStudy (anonymous):

 \[- \tan ^{2}x+\sec ^{2}x=-y ^{2}+\left( \sqrt{1+y ^{2}} \right)^{2}=-y ^{2}+1+y ^{2}=1\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

@surjithayer

OpenStudy (ozmosis):

leme make this a breeze for you, -tan^2(x)+sec^2(x)=1 we elaborate this further by expanding tan and sec \[-\frac{ \sin ^{2}x }{ \cos ^{2}x }+\frac{ 1 }{ \cos ^{2}x }=1\] we group the numerators above the already common denominator cos^2x \[\frac{ 1-\sin ^{2}x }{ \cos ^{2}x }=1\] we switch the numerators to get a better view of an already known trigonometric identity (1-sin^2x=cos^2x) \[\frac{ \cos ^{2}x }{ \cos ^{2}x }=1\] we divide cos^2x by itself, wich of course gives 1 \[1=1\]